Determine wheather the lines a parallel, perpendicular, or neither.
2x+5y=8
-5x+3=2y

Found 2 solutions by mananth, Lovely Jhea:
Answer by mananth(16946)

2x+5y=8 .............1
-5x-2y= -3 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 5
4x-y= 16
-25x+2y=-15
Add the two equations
-21x= 1
/-21
x=-0.05
plug value of x in (1) 2x+5y=8
-0.1+5y=8
5y=8+0.1
5y=8.1
y=1.62
They intersect at(0,1.62)
..................
CHECK by graphing
equation1
when x= 0,y=1.6(0,1.6)
when y =0,x=4(4,0)
Equation 2
when x= 0,y=1.5(0,1.5)
when y =0,x=0.6(0.6,0)




Answer by Lovely Jhea(1)

2+5=7 add 1 it will be 8
-5+3=-8 and -8 subtract by 2 is equal to 6 and 6 subtracted by 4 is equal to 2 then put y so then the answer is 2y :D